OpenStudy (anonymous):

what is the formula to find the circumference of a circle when the diameter is known??

OpenStudy (anonymous):

\(\large\color{black}{ \rm C=\pi d}\)

OpenStudy (anonymous):

C is the circumference, d is the diameter.

OpenStudy (anonymous):

C=πd

OpenStudy (anonymous):

you can use a 3.14 approximation for Pi.

OpenStudy (anonymous):

unless, you need an exact value.. whatever your teacher wants...

OpenStudy (anonymous):

c=2(3.14)r

OpenStudy (anonymous):

i just need to know the formula , and i had chose both of those, and they were wrong..

OpenStudy (anonymous):

\(\large\color{black}{ \rm C=\pi d}\) is the correct formula for the circumference (C), in terms of the diameter (d). For area on the other hand, you would have, \(\large\color{black}{ \rm A=\frac{\pi}{4}d}\)

OpenStudy (anonymous):

I mean d squared for the area... sorry

OpenStudy (anonymous):

would that be the same as pi times the radius squared ??